OpenStudy (anonymous):

Can someone help me solve this rational expression? 1/10 + 1/20 = 1/x. I need the steps, too.

OpenStudy (anonymous):

multiply through by l.c.m

OpenStudy (anonymous):

what would that be?

OpenStudy (anonymous):

lowest common multiple...

OpenStudy (anonymous):

which would be 20, so:\[\frac{ 1 }{ 10 } * 20 + \frac{ 1 }{ 20} *20 = \frac{ 1 }{ x } *20\]\[2 + 1 = \frac{ 20}{ x} \]Cross multiplying:\[3x = 20 \rightarrow x = \frac{ 20 }{ 3 }\]

OpenStudy (anonymous):

okay, that makes sense. thank you! I just wasn't sure what to do with the numerators. thank you so much!

OpenStudy (anonymous):

Yw :)